OpenStudy (anonymous):

Evaluate 20 ÷ (6 − 1) • 2

OpenStudy (jackellyn):

Remember PEMDAS! What would be your first step?

mathslover (mathslover):

20/5 * 2 = 4*2= 8

OpenStudy (anonymous):

first step is parenthesis

OpenStudy (jackellyn):

Awesome, then you're gonna either multiply or divide by what comes first, left to right. So what is next?

mathslover (mathslover):

according to VBODMAS rule we have to do division first .... hence 20/5 first and then . 5 = 4 * 2 = 8

OpenStudy (anonymous):

What does VBODMAS Stand for?

mathslover (mathslover):

V means BAR B means bracket O means of D means Division O means multiplication A means Addition S means Subtraction

OpenStudy (jackellyn):

is that number theory?

mathslover (mathslover):

yes ... this VBODMAS full form is just generally used in india and may be some other countries

OpenStudy (jackellyn):

it is similar to PEMDAS, which is how I learned it here in the US :)

mathslover (mathslover):

:)